A sample of gas in a balloon has an initial temperature of 36 ∘C and a volume of 1.17×103 L . If the temperature changes to 71 ∘C , and there is no change of pressure or amount of gas, what is the new volume, V2, of the gas?
Answer #1

Initial temperature, T1 = 36oC

Initial Volume = 1.17X103 L

Final temperature, T2 = 71oC

Final Volume, V2 = ?

We know that, at constant pressure and constant number of moles of gas,

V1 /T1 = V2 /T2

1.17X103 /36 = V2 / 71

V2 = 1.17X103 *71 / 36

V2 = 2.3075X103 L